I need help in with the following for questions. Thank you
1) 26 yds. 14 ft. 10 in - 13 yds 17 ft 12 in =
2) 10m i 17 yds 10 ft 5 in - 5 mi 95 yds 14 ft 16 in =
3) 12 ft 12 in - 11 ft 24 in =
4) 15 mi 28 ft - 6 mi 4,380 ft =

Answer by ilana(307)

I would think of this like regular subtraction. Suppose you wanted to find 543 - 365. You would start with ones. You can't do 3-5, so you borrow ten from the tens column. So you have 13-5, which is 8 ones. Now you look at tens. You borrowed one ten, so you are left with 3-6. You can't do that, so you borrow one from the hundreds column. So you have 13-6, or 7. Now you look at hundreds. You borrowed one hundred, so you are left with 4-3, or 1. So the answer is 1 hundred, 7 tens, and 8 ones, or 178.
1) You can't subtract 12in from 10in, so you borrow 1 foot. Since 1ft is 12 inches, you now have 22in-12in, or 10in. Now look at ft. You can't do 13ft-17ft, so borrow a yard. Since 1yd is 3ft, you now have 16ft-17ft, which you still can't do. So borrow another yd. So you have 19ft-17ft, or 2 ft. Now look at yds. You borrowed 2yd, so you have 24 left. 24yd-13yd = 11yd. So the answer is 11yd, 2ft, 10in.
Another way to complete these is to convert everything to the smallest unit, in this case inches, and then subtract. So you would have (936+168+10)in - (468+204+12)in = (1114-684)in = 430in. But then you probably need to convert that back to yds, ft, and inches, which is why I like the borrowing method better. Best of luck!